OpenStudy (anonymous):

What is the equation of a line that passes through the point (-2, 5) and is parallel to a line with a slope of 3?

OpenStudy (anonymous):

A. y = 3x + 1 B. y = 3x + 5 C. y = 3x + 7 D. y = 3x + 11

OpenStudy (mertsj):

You know a POINT on the line and the SLOPE of the line. Why don't you use POINT-SLOPE form?

OpenStudy (mertsj):

Which is: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (a1234):

Use point-slope form, y - y1 = m(x - x1) y - 5 = 3(x + 2) y - 5 = 3x + 6 y = 3x + 11 *Parallel lines have the same slope, but different y-intercepts.